OpenStudy (anonymous):

i got a few ugh 3+4(z+5)=31 5m+2(m+1)=23 27=3c-3(6-2c)

OpenStudy (texaschic101):

3 + 4(z + 5) = 31 3 + 4z + 20 = 31 4z = 31 - 20 - 3 4z = 31 - 23 4z = 8 z = 8/4 z = 2 ============ 5m + 2(m + 1) = 23 5m + 2m + 2 = 23 7m + 2 = 23 7m = 23 - 2 7m = 21 m = 21/7 m = 3 ============= 27 = 3c - 3(6 - 2c) 27 = 3c - 18 + 6c 27 + 18 = 3c + 6c 45 = 9c 45/9 = c 5 = c ==============